OpenStudy (anonymous):

Zach’s rectangular garden measure 12 ft by 10 ft. He puts a stake in each corner and one every 2 ft along each side. How many stakes are there in all?

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy! :) @kaykay_08 to do this, find the perimeter of the 12ft by 10ft rectangle so P = 12 + 12 + 10 + 10 = ? then since the stakes are every two feet, divide the number you just got by 2

OpenStudy (gorv):

that mean u need to calculate area

OpenStudy (gorv):

area=12*10

OpenStudy (gorv):

area of skate=2*2

OpenStudy (gorv):

no of skate=area of garden/area of skate